OpenStudy (anonymous):

Reversing the digits of Gyver's age gives Ed's age with a difference of 18 years. If the sum of the digits of each age is 6, how old is Gyver?

OpenStudy (butterflydreamer):

for this i think you should just start off with guess and check xD

OpenStudy (butterflydreamer):

so let's start off with two SINGLE DIGIT numbers that add up to give you 6.

OpenStudy (anonymous):

1 and 5, 4 and 2

OpenStudy (butterflydreamer):

yes so let's try 1 and 5 firstly. So the numbers we'd use would be 15 and 51. Now we need to find the difference to make sure it's = 18. So does 51 - 15 = 18 ?

OpenStudy (anonymous):

no :)

OpenStudy (butterflydreamer):

correct :) so that means 1 and 5 are not the numbers we'd use. So let's try 2 and 4. so our numbers would be 24 and 42 right? Does 42 - 24 = 18 ?

OpenStudy (anonymous):

yes :)

OpenStudy (butterflydreamer):

exactly ^_^. So now we have to figure out the age of Gyver

OpenStudy (butterflydreamer):

From the question we know that Gyver's age = Ed's age - 18 So if Gyver is 24, Ed would be 42 with a difference of 18 correct?

OpenStudy (anonymous):

oh so the answer is 24 :) thank you :)

OpenStudy (butterflydreamer):

i think it's 24 xD But now i'm confused if he is the one who is 42 or 24 :S ...

OpenStudy (butterflydreamer):

@Nnesha xD can you check? I can't work out if Gyver is the one that's 24 or 42

Nnesha (nnesha):

i never met him befr if i ever get a chance i'll ask

Nnesha (nnesha):

:P:D

Nnesha (nnesha):

i guess 24 is right :-)

OpenStudy (butterflydreamer):

omg xD you troll. Okey thanks for checking ^_^

OpenStudy (anonymous):

thank you again :)

OpenStudy (anonymous):

hi butterflydreamer do you thing you can help me with some math things

Nnesha (nnesha):

np :P

OpenStudy (anonymous):

Nnasha wow are you talking to?

OpenStudy (butterflydreamer):

no problem @martaamador62 :D (fingers crossed it is 24 not 42 xD) :) @Loveheart i can try

OpenStudy (anonymous):

wait how will i know if it's either 24 or 42 ? :)

OpenStudy (anonymous):

thanks for helping me :), i have an exit exam on math tom. :))

Nnesha (nnesha):

gO_Od luck!

OpenStudy (anonymous):

thanks @Nnesha :))

OpenStudy (butterflydreamer):

Well what i THINK it is issssss.. Gyver = 24 Then reverse it and we get ed's age 42. And they have a difference of 18 right? xD Since Ed - Gyver = 42 - 24 = 18 So yaaah... anyways GOOD LUCK :D !

OpenStudy (anonymous):

wait i think i manage to make an equation in order to know if it's 24 or 42 :)

OpenStudy (butterflydreamer):

oooOoOooo do share ^_^

OpenStudy (anonymous):

ohh thank you again :)

OpenStudy (butterflydreamer):

you're welcome times 3 xDD